using rule of 72, calculate you are hoping that your investment of$5000 will double in 7 years. what interest rate do you need to earn to make this happen
Answer #1

Investment Amount = PV = $5000

Let Interest Rate = r

Number of Years = 7

Future Value (FV) required after 7 years = 2*5000 = $10000

We know that FV = PV(1+r)n

=> 10000 = 5000(1+r)7

=> r = (10000/5000)1/7 - 1 = 0.1041 or 10.41%